OpenStudy (help_people):

multiply 3x^2(x^4-2x) and what exponent rule did you use?

OpenStudy (anonymous):

3x^6−6x^3

OpenStudy (anonymous):

Would you like to know how I got it ?

OpenStudy (help_people):

yes please and what exponent rule applies ?

OpenStudy (help_people):

@jenny1003

OpenStudy (anonymous):

=(3x2)(x4+−2x) =(3x2)(x4)+(3x2)(−2x) =3x6−6x3 I guess product rule but I could be wrong

OpenStudy (help_people):

ok :) ty could you help w. more at 5:00 ? @jenny1003

OpenStudy (help_people):

i have to do some stuff so...

OpenStudy (anonymous):

Ok jus @ me when you back on

OpenStudy (help_people):

wait may you explain to me what product rule is ?

OpenStudy (anonymous):

Rule (a n · a m = a n+m) Ex (23 · 24 = 23+4 = 128)

OpenStudy (help_people):

wait could you show with exponents (ik i am asking a lot)

OpenStudy (anonymous):

Sureee Lol

OpenStudy (anonymous):

oduct rule with same exponent an · bn = (a · b)n Example: 32 · 42 = (3·4)2 = 122 = 12·12 = 144

OpenStudy (help_people):

@jenny1003

OpenStudy (anonymous):

Do you get it or still confused?

OpenStudy (anonymous):

@help_people

OpenStudy (help_people):

I get it ty :) see you at 5 and when i do may you please explain everything ? lol i am just trying to understand and you promise you will be there?

OpenStudy (anonymous):

Like I said jus @ me and I will

OpenStudy (help_people):

and also could you do this if your showing an epxpontent ^ for example 2 squared = 2^2

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Are you asking what 2^2 is ?

OpenStudy (anonymous):

If so it's 4

OpenStudy (help_people):

nno i am saying that when you are using an exponent could you put the ^ sign Example 2 squared put 2^2

OpenStudy (help_people):

thats what I'm saying use the ^ sign cause i noticed you were not and it confused what was an exponent and what was not lol

OpenStudy (anonymous):

Lol yea you can

OpenStudy (anonymous):

I'm sorry

OpenStudy (help_people):

that ok but may you use it later on but i have to go in like a minute

OpenStudy (anonymous):

5 Right ? Jus @ me. See ya later

OpenStudy (help_people):

4:55 lol

OpenStudy (help_people):

but ok i will make a new tab bye then

OpenStudy (help_people):

* by

OpenStudy (anonymous):

Ok